A student takes a multiple choice test consisting of 5 questions where there are 4 choices per question. Also, each question is worth 20 points. Suppose the student guesses on each question. Let X be the number of questions the student gets correct. The probability distribution for X is given below. Find the probability that the student gets at least a 60 on the test. Give your answer to four decimal places.
Answer #1

Given problem under Binomial distribution

p=\frac{1}{4}=0.25 Let q=1-p=0.75 and n=5

To find P(X\geq 3)

P(X\geq 3)=P(X=3)+P(X=4)+P(X=5)

  =5_{C_{3}}(0.75)^{5-3}(0.25)^{3}+5_{C_{4}}(0.75)^{5-4}(0.25)^{4} +5_{C_{5}}(0.75)^{5-5}(0.25)^{5}

=10(0.75)^{2}(0.25)^{3}+5(0.75)(0.25)^{4}+(0.25)^{5}

  =0.1035